# IoT Ground Sensing Systems for Early Wildfire Detection: Technologies, Challenges and Opportunities

### v1

@@ II-B1 Weather Monitoring Sensor
The utilisation of temperature, humidity and air pressure detection technologies is prevalent in IoT projects, with thermistors and resistive temperature detectors (RTDs) commonly used for temperature sensing [52, 53]. Similarly, humidity and air pressure sensing are employed for resistive humidity sensing, thermal conductivity humidity sensing and piezoresistive pressure sensing, respectively, considering factors such as size and cost in mass production. Weather sensors have the potential to be an important asset in monitoring fire risk, as temperature and relative humidity can be used to calculate the moisture content of fire fuels, which is a major factor in controlling fire activity [54], including ignition probability [55]. Furthermore, it allows the detection of indicators such as increased temperature, humidity reduction and air pressure changes resulting from fire heat, which can be observed by EDs. These indicators are commonly used to determine the spread of wildfires, as well as are widely used in various indoor fire detection schemes [56, 57, 58].
@@
Thermistors exhibit high sensitivity to temperature changes as a result of the heat conductivity of their sensing materials, resulting in a lower negative temperature coefficient and a nonlinear response. However, thermistors have limited operating temperature ranges and may not be optimal as standalone temperature sensors for IoT fire detection. On the other hand, RTDs offer wider temperature operating ranges but have lower sensitivity compared to thermistors. In humidity detection, both capacitive and resistive humidity sensors are cost-effective. However, capacitive humidity sensors are sensitive to contaminants and temperature variations, while resistive humidity sensors are generally larger in size, which can impact the design considerations of IoT projects. [59, 60] have demonstrated the potential application of these sensing technologies to detect fire hazards in forested areas. A recent study [61] comprehensively analysed the accuracy of several available temperature and humidity sensors on the market. The findings revealed that BME280 exhibited the highest precision, even under extreme weather conditions. However, AM2302E showed higher errors at low temperatures, while AM2302F and SHT71 had a higher error rate within the 30-60 humidity range. Although its successor, AM2321, demonstrated a slight improvement, it remains sensitive to temperature changes. BME280 was considered the most accurate. However, relying solely on weather sensors may prove insufficient for timely wildfire detection due to their slow rate of change.
@@
To assess the responsiveness and effectiveness of weather sensors in detecting fire incidents, our team carried out an experimental burn in a national park. In this experiment, we placed multiple BME280 sensors 10 metres from the source of the fire, evenly spaced in a circle. Fig. 3 presents the average temperature and relative humidity data obtained from the sensors at each time. We ignited two fires using approximately 1.6 kg of Eucalyptus leaves (equivalent to 22.535 tonnes per hectare) in a 0.7 square metered fire pit. The fires were lit around 11:00 and 12:00, each lasting ten minutes with an additional five minutes of smouldering, but we can see from Fig. 3 that the weather sensors did not identify any notable irregularities. It suggests that depending only on weather sensors may not be sufficient to identify fires in their initial stages, particularly when the same experiment configuration is tested for gas sensors at a distance of 30 metres from the fire sources, as demonstrated in Fig. 4. However, the fires used in the example represent the lowest limit of detectable fire size. Fires up to 5 hectares can still be suppressed under mild weather conditions [5], and may alter the local temperature and humidity more than the fire in this example. Besides, as mentioned above in this subsection, fuel moisture, which is important for ignition success, can still be calculated from temperature, relative humidity and air pressure [54, 55]. Therefore, the moisture content of the vegetation can serve as the wildfire risk indicator and provide valuable guidance on the frequency of sensor readings required for early wildfire detection.
@@ II-B2 Gas / Smoke Sensors
During the early stages of a wildfire, the smouldering process can result in the release of a significant amount of greenhouse gases and heat. The gases emitted by a small fire may disperse before forming a visible plume [62]. Although heat can be detected by humidity, thermal and pressure sensors, a large amount of gases, such as carbon dioxide (), methane () and nitrous oxide () [63], together with particulate matter (PM), are also released from wildfires and can be measured using gas/smoke sensors. Previous research has shown that most of the carbon emitted during wildfires is in the form of (88), with smaller amounts of carbon monoxide (6) and organic compounds without (3.8) [64]. As a result, gas/smoke sensors capable of detecting concentrations of these gases or smoke particles are often used in wildfire detection systems. Various types of sensors, including metal oxide semiconductors, chemical field-effect transistors and electrochemical sensors, are currently used to measure the concentrations of total volatile organic compounds (TVOC) and CO in the atmosphere, while optical or laser scattering sensors are commonly used to detect PM particles. These sensors provide gas concentration readings at a high sampling rate to MCUs or to the cloud server, which can facilitate early fire detection. In particular, sudden changes in gas concentrations are often more detectable than gradual changes in environmental factors such as temperature and humidity, making wildfire detection easier using such gas/smoke concentration data. We conducted another experiment with the same parameters as the one shown in Fig. 3, with sensors placed 30 metres from the fire and with multiple ignitions. Fig. 4 illustrates a strong connection between fires, CO and TVOC levels, while PM concentration is strongly associated with wildfires, as demonstrated in [65].
@@
Despite the availability of these sensors on the market, it is still necessary to compare their characteristics and performance. Therefore, further research efforts are necessary to determine the most suitable sensor type for an IoT-based wildfire detection project. In the following sections, we present a review of various sensor types that exhibit the potential to detect smoke PM, CO and TVOC gases.
@@
Resistive Metal Oxide Semiconductor Sensor (MoS): MoS sensors have gained widespread usage in a variety of IoT projects because of their cost-effectiveness and popularity in IoT development. The operational principle of MoS sensors involves the controlled heating of a metal oxide plate to a specific temperature, typically within the range of 300-400C. This temperature enables the metal oxide to react with the target gas, leading to a chemical reaction that reduces the electrical resistance of the semiconductor. Therefore, changes in the resistance of the MoS sensor can be used to determine the concentration of the target gas. These sensors are particularly well suited for early stage wildfire detection systems in outdoor environments, because of their affordability and ability to be deployed on a large scale. Metal oxide gas sensors exhibit robustness and stability in outdoor settings, allowing them to withstand various weather conditions, temperature fluctuations and humidity levels. As a result, they offer reliable operation for extended periods in forested areas [66].
@@
Despite the advantages of MoS sensors, they are known for their relatively high energy consumption, as they require energy to heat the metal oxide plates. Additionally, extreme weather conditions, such as high humidity, can adversely affect their performance. Therefore, protective measures, such as placing sensors in weatherproof enclosures, are necessary in such conditions. MoS sensors have a high sensitivity to a wide range of gases, allowing for a quick response to the target gas, but this heightened sensitivity also introduces a high cross-sensitivity to other gases, leading to inaccurate readings, particularly in the presence of smoke. Furthermore, MoS sensors need a period of time to reach the temperature needed for precise measurements. As a result, the baseline reading of the MoS sensors can change over time due to variations in temperature, changes in humidity and exposure to different gases, as illustrated in Fig. 5. Consequently, periodic calibration of the baseline is necessary when operating for extended periods.
@@
In [72], the response time and the recovery time of various metal oxide catalysts were evaluated at a distance of 6 cm from a concentration of 500 ppm of SO and the minimum response time was found to be 80 seconds, while the minimum recovery time was 70 seconds accordingly. Catalytic sensors are more accurate than MoS sensors in detecting combustible gases such as Methane, Ethane, Butane and other VOCs. This is because catalytic sensors have low cross-sensitivity, meaning that they are less likely to respond to other gases. As a result, catalytic sensors require less calibration effort than MoS sensors.
@@
Similarly to MoS sensors, the utilisation of catalytic beads requires elevated temperatures to activate the catalytic pellets, resulting in shorter life and higher energy consumption. Furthermore, it should be noted that for IoT systems in wildfire scenarios, where most released gases are composed of CO, catalytic sensors may not be the most suitable choice for detection, as they are designed primarily for combustible gases. Furthermore, catalytic sensors are highly sensitive to temperature fluctuations and tend to be more expensive than MoS sensors. They are also prone to contamination when exposed to high concentrations of combustible gases, which can lead to degradation in sensor performance. Therefore, catalytic beads sensors are used primarily in indoor environments, such as process plants and cargo compartments of aircraft, due to their suitability for these specific applications [73, 74].
@@
Electro-Chemical Sensor: Electrochemical sensors have been widely used in indoor fire detection and environmental monitoring projects in the past decade. However, because of its working principle, the miniaturisation of the electrochemical sensor becomes very limited and becomes less popular than the MoS sensor technology for early wildfire detection projects. Typical electrochemical sensors comprise an electrode that interfaces with an electrolyte solution, enabling the detection of target gases. When the gas is exposed, it diffuses through a porous membrane surrounding the electrode, initiating oxidisation or reduction reactions, and generating an electric current within the circuit. The gas concentration is then determined through the measured current. Compared to MoS sensors, electrochemical sensors exhibit higher selectivity toward CO and CO due to their use of chemical reactions for gas detection, while MoS sensors are based on the physical properties of the material. Direct measurement of current induced by chemical reactions in electrochemical sensors offers selectivity higher than that of the physical properties exploited by MoS sensors, making the former more effective in detecting specific gases. Furthermore, electrochemical sensors generally have longer lifetimes and lower power consumption because their electrodes do not require heating to allow chemical reaction. However, their operating temperature range can be limited, and their sensing range may be restricted.
@@
Although several researchers have dedicated their efforts to advancing these sensor technologies [81, 83, 84], it is challenging to find FET-type sensors readily available on the market for the detection of gases such as CO, CO or TVOCs. Ongoing research and development efforts focus on the advancement of FET-type gas sensors with substantial potential for their application in various fields. The future holds promising prospects for the wide availability and utilisation of FET-type gas sensors in diverse applications.
@@
The performance of gas/smoke sensors can vary greatly, making it difficult for IoT designers to understand the difference in each sensor performance. To address this, many research efforts have been made to compare the accuracy, response time, selectivity and cross-sensitivity of sensors. A comparative study conducted by [87] examined the sensitivity and accuracy of MoS, GasFET, Electro-Chemical, NDIR and optical scattering detectors. The results showed that the MoS and GasFET sensors had sensitivity comparable to those of the other sensor types. Furthermore, [57] provided an extensive review of recent developments in fire detection sensors, including an explanation of their operating principles and a comparison of the performance of different types of sensors using different detection elements.
@@
To help IoT designers understand the features of these gas/smoke detectors, we have compiled a comparison of the performance of various types of gas/smoke sensors that are suitable for IoT projects for early wildfire detection in Table II. This analysis is based on commercially available sensors and takes into account factors such as cost, energy usage, reaction time, durability and precision. These aspects are beneficial to the IoT designer and can help them select the right sensor for their project. In addition, the most remarkable characteristics of the sensors are also included in the table. Furthermore, Table III presents an overview of various IoT-based wildfire detection initiatives and the types of ground sensors used in each. It shows that many wildfire detection projects rely mainly on environmental monitoring sensors, including both weather monitoring and gas/smoke sensors, and MoS sensors are the most widely used gas/smoke sensors in these projects. In [80, 88], a comprehensive solution is presented, from hardware design to detection flow and DL detection, while [89] also offers a complete solution with implemented ML detection. [90] provides an experimental hardware prototype, but also includes a fire-carbon spreading model which can help determine the optimal detection distance and deployment height. [91] presents a real project of the deployed wildfire detection system in the United States and also provides information on the sensor topology. [38, 70, 92, 93, 94, 95] present experimental proof-of-concepts of IoT ground sensing systems and different measurements are taken using their setup, while [93] demonstrates the feasibility of mounting environmental monitoring sensors on UAVs to detect early wildfires. Although [96] is only an example of the real deployment of the PM2.5 IoT sensing system, it also shows the ability of on board ML detection and the robustness of the use of LoRaWAN connectivity. [31, 29, 97] are real commercial IoT projects for early wildfire detection. Lastly, [98] provides an IoT system for early fire detection featuring GaSFET sensors.

### v2

@@ II-B1 Weather Monitoring Sensor
The utilisation of temperature, humidity and air pressure detection technologies is prevalent in IoT projects, with thermistors and resistive temperature detectors (RTDs) commonly used for temperature sensing [53, 54]. Similarly, humidity and air pressure sensing are employed for resistive humidity sensing, thermal conductivity humidity sensing and piezoresistive pressure sensing, respectively, considering factors such as size and cost in mass production. Weather sensors have the potential to be an important asset in monitoring fire risk, as temperature and relative humidity can be used to calculate the moisture content of fire fuels, which is a major factor in controlling fire activity [55], including ignition probability [56]. Furthermore, it allows the detection of indicators such as increased temperature, humidity reduction and air pressure changes resulting from fire heat, which can be observed by EDs. These indicators are commonly used to determine the spread of wildfires, as well as widely used in various indoor fire detection schemes [57, 58].
@@
Thermistors exhibit high sensitivity to temperature changes as a result of the heat conductivity of their sensing materials, resulting in a lower negative temperature coefficient and a nonlinear response. However, thermistors have limited operating temperature ranges and may not be optimal as standalone temperature sensors for IoT fire detection. On the other hand, RTDs offer wider temperature operating ranges but have lower sensitivity compared to thermistors. In humidity detection, both capacitive and resistive humidity sensors are cost-effective. However, capacitive humidity sensors are sensitive to contaminants and temperature variations, while resistive humidity sensors are generally larger in size, which can impact the design considerations of IoT projects. [59, 60] have demonstrated the potential application of these sensing technologies to detect fire hazards in forested areas. A recent study [61] comprehensively analysed the accuracy of several available temperature and humidity sensors on the market. The findings revealed that BME280 exhibited the highest precision, even under extreme weather conditions. AM2302E showed higher errors at low temperatures, while AM2302F and SHT71 had a higher error rate within the 30-60 humidity range. Although its successor, AM2321, demonstrated a slight improvement, it remains sensitive to temperature changes. BME280 was considered the most accurate. The BME680 is the successor to the BME280, distinguished by its ability to detect TVOCs. Both sensors demonstrate similar accuracy in their measurements, with consistency in humidity, barometric pressure, and temperature. However, the BME680 exhibits a marginally higher humidity hysteresis of 1.5, in contrast to the 1.2 observed in the BME280. Furthermore, the BME680 has a reduced root mean square noise in pressure measurements, at 0.12 hPa, compared to the 0.2 hPa of the BME280 [62, 63]. Nevertheless, relying solely on weather sensors may prove insufficient for timely wildfire detection due to their slow rate of change.
@@
In order to evaluate the ability of weather sensors in detecting fire, our team conducted several experimental burns in a national park on various days during the low-risk wildfire season. In this set of experiments, we placed several BME280 sensors evenly spaced in a circle at varying distances, ranging from 10 to 50 meters, from the fire source. It should be noted that the BME280 sensors are highly accurate micro-sensors that can be used in IoT devices, as previously mentioned. These sensors are also used in various electronic devices, such as smart watches, smart home monitors, and other wildfire detection systems. The temperature and relative humidity changes obtained from the sensors during a typical experimental burn are shown in Fig. 3. The sensors were positioned 10 metres away from the fire source. In this experimental burn, we ignited two fires using approximately 1.6 kg of Eucalyptus leaves (equivalent to 22.535 tonnes per hectare) in a 0.7 square metered fire pit. The fires were lit at around 60 and 120 minutes after activating the sensors, each lasting ten minutes with an additional five minutes of smouldering, but we can see from Fig. 3 that the weather sensors did not identify any notable irregularities. It suggests that depending only on weather sensors may not be sufficient to identify fires in their initial stages, particularly when the same experiment configuration is tested for gas sensors at a distance of 30 metres from the fire sources, as demonstrated in Fig. 4. However, the fires used in the example represent the lowest limit of detectable fire size. Fires up to 5 hectares can still be suppressed under mild weather conditions [5], and may alter the local temperature and humidity more than the fire in this example. Besides, as mentioned above in this subsection, fuel moisture, which is important for ignition success, can still be calculated from temperature, relative humidity and air pressure [55, 56]. Therefore, the moisture content of the vegetation can serve as the wildfire risk indicator and provide valuable guidance on the frequency of sensor readings required for early wildfire detection.
@@ II-B2 Gas / Smoke Sensors
During the early stages of a wildfire, the smouldering process can result in the release of a significant amount of greenhouse gases and heat. The gases emitted by a small fire may disperse before forming a visible plume [64]. Although heat can be detected by humidity, thermal and pressure sensors, a large amount of gases, such as carbon dioxide (), methane () and nitrous oxide () [65], together with particulate matter (PM), are also released from wildfires and can be measured using gas/smoke sensors. Previous research has shown that most of the carbon emitted during wildfires is in the form of (88), with smaller amounts of carbon monoxide (6) and organic compounds without (3.8) [66]. As a result, gas/smoke sensors capable of detecting concentrations of these gases or smoke particles are often used in wildfire detection systems.
Various types of sensors, including metal oxide semiconductors, chemical field-effect transistors and electrochemical sensors, are currently used to measure the concentrations of total volatile organic compounds (TVOC) and CO2 in the atmosphere, while optical or laser scattering sensors are commonly used to detect PM particles. These sensors provide gas concentration readings at a high sampling rate to MCUs or to the cloud server, which can facilitate early fire detection. In particular, sudden changes in gas concentrations are often more detectable than gradual changes in environmental factors such as temperature and humidity, making wildfire detection easier using such gas/smoke concentration data. We conducted another experiment with the same parameters as the one shown in Fig. 3, with sensors placed 30 metres from the fire and with multiple ignitions. Figure 4 shows a clear correlation between fires, levels of CO2, and TVOCs. The graph shows that there are multiple peaks in TVOC and CO2 concentrations during the burning periods, indicating that the change in these gas concentrations due to wildfires is much more substantial than for wildfires compared to other meteorological factors such as temperature and humidity.
@@
Despite the availability of these sensors on the market, it is still necessary to compare their characteristics and performance. Therefore, further research efforts are necessary to determine the most suitable sensor type for an IoT-based wildfire detection project. In the following sections, we present a survey of various sensor types that exhibit the potential to detect smoke PM, CO2 and TVOC gases.
@@
Resistive Metal Oxide Semiconductor Sensor (MoS): MoS sensors have gained widespread usage in a variety of IoT projects because of their cost-effectiveness and popularity in IoT development. The operational principle of MoS sensors involves the controlled heating of a metal oxide plate to a specific temperature, typically within the range of 300-400∘C. This temperature enables the metal oxide to react with the target gas, leading to a chemical reaction that reduces the electrical resistance of the semiconductor. Therefore, changes in the resistance of the MoS sensor can be used to determine the concentration of the target gas. These sensors are particularly well suited for early stage wildfire detection systems in outdoor environments, because of their affordability and ability to be deployed on a large scale. Metal oxide gas sensors exhibit robustness and stability in outdoor settings, allowing them to withstand various weather conditions, temperature fluctuations and humidity levels. As a result, they offer reliable operation for extended periods in forested areas [67].
@@
Despite the advantages of MoS sensors, they are known for their relatively high energy consumption, as they require energy to heat the metal oxide plates. Additionally, extreme weather conditions, such as high humidity, can adversely affect their performance. Therefore, protective measures, such as placing sensors in weatherproof enclosures, are necessary in such conditions. MoS sensors have a high sensitivity to a wide range of gases, allowing for a quick response to the target gas, but this heightened sensitivity also introduces a high cross-sensitivity to other gases, leading to inaccurate readings, particularly in the presence of smoke. Furthermore, MoS sensors need a period of time to reach the temperature needed for precise measurements. As a result, the baseline reading of the MoS sensors can change over time due to variations in temperature, changes in humidity, and exposure to different gases, as illustrated in Fig. 5. Consequently, periodic calibration of the baseline is necessary when operating for extended periods. Note that other MQ series are also designed to detect various flammable gases. In this context, we have chosen the MQ2 sensor due to its wide detection range for gases emitted by wildfires, as previously mentioned in this subsection.
@@
In [72], the response time and the recovery time of various metal oxide catalysts were evaluated at a distance of 6 cm from a concentration of 500 ppm of SO2 and the minimum response time was found to be 80 seconds, while the minimum recovery time was 70 seconds accordingly. Catalytic sensors are more accurate than MoS sensors in detecting combustible gases such as Methane, Ethane, Butane and other VOCs. This is because catalytic sensors have low cross-sensitivity, meaning that they are less likely to respond to other gases. As a result, catalytic sensors require less calibration effort than MoS sensors.
@@
Similarly to MoS sensors, the utilisation of catalytic beads requires elevated temperatures to activate the catalytic pellets, resulting in shorter life and higher energy consumption. Furthermore, it should be noted that for IoT systems in wildfire scenarios, where most released gases are composed of CO2, catalytic sensors may not be the most suitable choice for detection, as they are designed primarily for combustible gases. Furthermore, catalytic sensors are highly sensitive to temperature fluctuations and tend to be more expensive than MoS sensors. They are also prone to contamination when exposed to high concentrations of combustible gases, which can lead to degradation in sensor performance. Therefore, catalytic beads sensors are used primarily in indoor environments, such as process plants and cargo compartments of aircraft, due to their suitability for these specific applications [73, 74].
@@
Electro-Chemical Sensor: Electrochemical sensors have been widely used in indoor fire detection and environmental monitoring projects in the past decade. However, because of its working principle, the miniaturisation of the electrochemical sensor becomes very limited and becomes less popular than the MoS sensor technology for early wildfire detection projects. Typical electrochemical sensors comprise an electrode that interfaces with an electrolyte solution, enabling the detection of target gases. When the gas is exposed, it diffuses through a porous membrane surrounding the electrode, initiating oxidisation or reduction reactions, and generating an electric current within the circuit. The gas concentration is then determined through the measured current. Compared to MoS sensors, electrochemical sensors exhibit higher selectivity toward CO2 and CO due to their use of chemical reactions for gas detection, while MoS sensors are based on the physical properties of the material. Direct measurement of current induced by chemical reactions in electrochemical sensors offers selectivity higher than that of the physical properties exploited by MoS sensors, making the former more effective in detecting specific gases. Furthermore, electrochemical sensors generally have longer lifetimes and lower power consumption because their electrodes do not require heating to allow chemical reaction. However, their operating temperature range can be limited, and their sensing range may be restricted.
@@
Although several researchers have dedicated their efforts to advancing these sensor technologies [82, 84, 85], it is challenging to find FET-type sensors readily available on the market for the detection of gases such as CO, CO2 or TVOCs. Ongoing research and development efforts focus on the advancement of FET-type gas sensors with substantial potential for their application in various fields. The future holds promising prospects for the wide availability and utilisation of FET-type gas sensors in diverse applications.
@@
The performance of gas/smoke sensors can vary greatly, making it difficult for IoT designers to understand the difference in each sensor performance. To address this, many research efforts have been made to compare the accuracy, response time, selectivity and cross-sensitivity of sensors. A comparative study conducted by [88] examined the sensitivity and accuracy of MoS, GasFET, Electro-Chemical, NDIR and optical scattering detectors. The results showed that the MoS and GasFET sensors had a sensitivity comparable to that of the other sensor types, while MoS, FET, catalytic, and electrochemical sensors display a faster response time, whereas NDIR sensors are typically slower by about 15 seconds compared to others. This observation has been further investigated by the same group in an obstructed scenario, where it was observed that the response time of the NDIR sensors could be delayed further due to blockages of the heat layers [89]. In addition, comparisons among various MoS sensors have been documented [90, 91]. SGP30 has been observed to typically react to most VOCs faster than BME688 and ENS160, with a difference of 3–15 seconds, while ENS160 tends to produce more noisy readings. Furthermore, [58] provided an extensive review of recent developments in fire detection sensors, including an explanation of their operating principles and a comparison of the performance of different types of sensors using different detection elements.
@@
To help IoT designers understand the features of these gas/smoke detectors, we have compiled a comparison of the performance of various types of gas/smoke sensors that are suitable for IoT projects for early wildfire detection in Table III. This analysis is based on commercially available sensors and takes into account factors such as cost, energy usage, reaction time, durability and precision. These aspects are beneficial to the IoT designer and can help them select the right sensor for their project. In addition, the most remarkable characteristics of the sensors are also included in the table. Furthermore, Table IV presents an overview of various IoT-based wildfire detection initiatives and the types of ground sensors used in each. It shows that many wildfire detection projects rely mainly on environmental monitoring sensors, including both weather monitoring and gas/smoke sensors, and MoS sensors are the most widely used gas/smoke sensors in these projects. In [81, 92], a comprehensive solution is presented, from hardware design to detection flow and DL detection, while [93] also offers a complete solution with implemented ML detection. [94] provides an experimental hardware prototype, but also includes a fire-carbon spreading model which can help determine the optimal detection distance and deployment height. [95] presents a real project of the deployed wildfire detection system in the United States and also provides information on the sensor topology. [42, 70, 96, 97, 98, 99] present experimental proof-of-concepts of IoT ground sensing systems and different measurements are taken using their setup, while [97] demonstrates the feasibility of mounting environmental monitoring sensors on UAVs to detect early wildfires. Although [100] is only an example of the real deployment of the PM2.5 IoT sensing system, it also shows the ability of on board ML detection and the robustness of the use of LoRaWAN connectivity. [39, 37, 101] are real commercial IoT projects for early wildfire detection. Additionaly, [102] provides an IoT system for early fire detection featuring GaSFET sensors. In [95, 42], both vision-based and environmental monitoring sensors were implemented on the same ED, but this approach cannot optimise the use of both technologies. To address this, [103] proposed a hierarchical sensor network featuring separate camera end nodes and sensor end nodes. This network employs ground sensors to detect temperature and humidity anomalies, subsequently confirming fires using a CNN camera imaging model. Although this method may result in a relatively longer detection time compared to using vision-based or environmental sensing-based detection alone, it has demonstrated an improve accuracy to 99. Furthermore, IoT sensors assist in filtering fire alarms, allowing the camera module to remain in sleep mode most of the time, thus significantly reducing the energy consumption from camera equipped EDs.